# Solving the Bongard-logo Problem by Modeling a Probabilistic Model

### v3

@@ IV Pose-tansformer
Previous research has demonstrated limited success in learning abstract concepts related to positional relationships among entities, particularly when addressing the Bongard-Logo and RPM problems. This limitation is evident in challenges posed by the OIG and D-9 problem sets from RAVEN, as well as the ff and hd issues of Bongard-Logo and the Neurtal problem in PGM. The RS-model, however, overcame these difficulties by explicitly incorporating additional structural cues (meta data, auxiliary label), achieving commendable inferential accuracy.
@@
In this paper, we introduce the Pose-Transformer, an enhancement to the Transformer-Encoder framework that incorporates the pose matrix from the Capsule Network[33]. This integration implicitly enhances the network’s ability to discern positional relationships among local representations, thereby bolstering its learning capabilities on graphical abstract reasoning problem. We present and delve into the intricacies of our proposed Pose-Transformer, discussing its potential to revolutionize learning in tasks that involve complex positional relationships through experiments.
@@
V Straw Pose-tansformer
@@
The pose vectors $\{p_{jk}|j\in[1,m],k\in[1,n],p_{jk}\in R^{d}\}$ is achieved by multiplying the attention heads with the mapping matrices. These pose vectors, after summation and squashing, are added to the vector set $\{V_{i}|i\in[1,n],V_{i}\in R^{d}\}$ to obtain the output of new pose matrix mapping block $\{O_{i}|i\in[1,n],O_{i}\in R^{d}\}$. This process outlines the computation process of the pose embedding matrix block within the Straw Pose-Transformer architecture. The entire calculation process of new “pose matrix embedding” can be expressed as fallows:
@@
In the formula, the process of the Masked transformer-encoder to obtain the output is denoted as “Masked T-Encoder”. In the Straw Pose-Transformer architecture, the outputs from each logical layer, denoted as $\{O_{k}|k\in[1,n]\}$, undergo a concatenation process with the global vector $V_{0}$ attached to their precursors. This modified set, now represented as $\{\tilde{V}_{i}|i\in[0,n]\}$ , is then propagated into the subsequent logical layer for further processing. However, if the subsequent logical layer is the final layer, the global vector $V_{0}$ is no longer appended.
In this way, we can construct the network according to the formula, and the feedforward process of a single logical layer is illustrated in the figure 10. It is worth mentioning that masking the position corresponding to $V_{0}$ in the attention coefficients reduces the impact of the optimizable parameter $V_{0}$ on the pose mapping process to a certain extent, making it a relatively necessary operation. our analysis demonstrates that this stabilizes training and improves inference accuracy somewhat. This paper replaces the standard Transformer-Encoder+cross-attention+pose matrix mapping sequential process in pose-Transformer with the aforementioned procedure. And figure 11 illustrate the structure of the Straw-Pose-Transformer.
@@
Figure 11: The architecture of the straw pose-Transformer
@@
VI combination of Pose-Transformer and baselines
@@ VII Experiment
VII-A Experiment on Bongard-Logo
We conducted experiments on the Bongard-logo dataset using the SBSD and PMoC model. The results are recorded in Table I. Due to the serial convolution scanning mechanism and attention mechanism being prone to attention collapse, the parameters of the SBDC model are utilized as pre-trained parameters for the PMoC convolutional layer. In addition, this study experimented with the combination of the Straw pose-transformer and PMoC. It is worth noting that, under the experimental conditions of this paper, which involved four A100s graphics processing units (GPUs), the integration of PMoC with the pose-transformer proved to be challenging to deploy. This is one of the reasons and motivations behind our decision to design a lightweight approach for Pose-Transformer in this paper.
@@ VII-A Experiment on Bongard-Logo
TABLE I: Reasoning Accuracies of DC on Bongard-logo.
@@
PMoC 92.0 92.6 97.7 78.3 75.0
PMoC+Straw Pose-Transformer 96.2 96.0 98.5 78.5 75.3
In Table I, it is evident that addressing the Bongard-Logo problem using a probabilistic model approach, namely PMoC, is both straightforward and effective. Furthermore, SBSD exhibits stronger outcomes when compared to previous models. Additionally, the integration of Straw-Pose-Transformer has contributed significantly to enhancing the inference accuracy of PMoC.
VII-B Experiment on RPM
We conducted experiments on I-RAVEN and PGM by sequentially replacing the Transformer structure in RS-Tran with Pose-Transformer and Straw-Pose-Transformer. The results of I-RAVEN are recorded in Table II and PGM is recorded in Table III. In this paper, we considered that compared to the standard Transformer-Encoder, a single layer of Pose-Transformer and its lightweight version possess twice the number of logical computations. Therefore, the combined RS-Pose-Tran and RS-Straw-Pose-Tran only stacked half of the logical layers of RS-Tran to ensure that the internal logical computation times of the three models were equivalent. In addition, the parameter settings and dataset configurations for the experiments related to RS-Pose-Tran were strictly maintained consistent with those of RS-Tran. Such a setting renders the comparative results of the three models more convincing.
TABLE II: Reasoning Accuracies on RAVEN and I-RAVEN.
@@ VII-B Experiment on RPM
In Table III, it can be seen that we did not conduct experiments on RS-Pose-Tran for PGM. This is because compared to RS-Tran operating on the RAVEN dataset, the overall dimension of the RS-Tran model doubles on PGM. However, such a doubling is prohibitive for RS-Pose-Tran, making it undeployable on our equipment and device. This motivated us to lightweight the Pose-Transformer into the Straw-Pose-Transformer in this paper.
TABLE III: Reasoning Accuracies on PGM.
@@
RS-Straw-Pose-TRAN 98.64
From the perspective of experimental results, it is not difficult to find that our lightweight approach has been successful without significantly reducing the performance of RS-Pose-Tran. Furthermore, the introduction of the pose mapping matrix into the Transformer-Encoder is effective, significantly improving the performance of RS-Tran. This demonstrates that Pose-Transformer is more suitable for abstract reasoning capabilities compared to the standard Transformer-Encoder, even though the standard Transformer-Encoder structure has shown strong advantages in RPM and Bongard-logo problems.
VIII Conclusion
@@
In conclusion, our study highlights the effectiveness of incorporating pose mapping matrix into Transformer-Encoder for improving abstract reasoning capabilities. The successful implementation of our lightweight approach without significant performance reduction further strengthens the potential of Pose-Transformer in addressing complex reasoning tasks. The ingenious design of PMoC avoids to directly constrain the distribution of the latent representation of Bongard-Logo images, as well as the rigid modeling of this distribution into a known form. Experimental results demonstrate that such a design choice in PMoC is highly reasonable and effective. Furthermore, Pose-Transformer, leveraging PMoC, has exhibited its superiority in addressing the Bongard-logo problem.

### v7

@@ IV Methodology for abstract reasoning
Previous research has demonstrated limited success in learning abstract concepts related to positional relationships among entities, particularly when addressing the Bongard-Logo and RPM problems. This limitation is evident in challenges posed by the OIG and D-9 problem sets from RAVEN, as well as the ff and hd issues of Bongard-Logo and the Neurtal problem in PGM. The RS-model[31], however, overcame these difficulties by explicitly incorporating additional structural cues (meta data, auxiliary label), achieving commendable inferential accuracy.
@@
In this section, we introduce the Pose-Transformer, an enhancement to the Transformer-Encoder framework that incorporates the pose matrix from the Capsule Network[33]. This integration implicitly enhances the network’s ability to discern positional relationships among local representations, thereby bolstering its learning capabilities on graphical abstract reasoning problem. Furthermore, we have lightweighted Pose-Transformer into Straw-Pose Transformer, which achieves a lightweight version of the former while maintaining its effectiveness. We present and delve into the intricacies of our proposed Pose-Transformer, discussing its potential to revolutionize learning in tasks that involve complex positional relationships through experiments.
IV-A Pose-tansformer
@@
IV-B Straw Pose-Tansformer
@@
The pose vectors $\{p_{jk}|j\in[1,m],k\in[1,n],p_{jk}\in R^{d}\}$ is achieved by multiplying the attention heads with the mapping matrices. These pose vectors, after summation and squashing, are added to the vector set $\{V_{i}|i\in[1,n],V_{i}\in R^{d}\}$ to obtain the output of new pose matrix mapping block $\{O_{i}|i\in[1,n],O_{i}\in R^{d}\}$. This process outlines the computation process of the pose embedding matrix block within the Straw-Pose-Transformer architecture. The entire calculation process of new “pose matrix embedding” can be expressed as fallows:
@@
In the formula, the process of the Masked transformer-encoder to obtain the output is denoted as “Masked T-Encoder”. In the Straw-Pose-Transformer architecture, the outputs from each logical layer, denoted as $\{O_{k}|k\in[1,n]\}$, undergo a concatenation process with the global vector $V_{0}$ attached to their precursors. This modified set, now represented as $\{\tilde{V}_{i}|i\in[0,n]\}$ , is then propagated into the subsequent logical layer for further processing. However, if the subsequent logical layer is the final layer, the global vector $V_{0}$ is no longer appended.
In this way, we can reconstruct the pose matrix embedding process according to the formula, and this feedforward process is illustrated in the figure 10. It is worth mentioning that masking the position corresponding to $V_{0}$ in the attention coefficients reduces the impact of the optimizable parameter $V_{0}$ on the pose mapping process to a certain extent, making it a relatively necessary operation. our analysis demonstrates that this stabilizes training and improves inference accuracy somewhat. This paper replaces the standard Transformer-Encoder+cross-attention+pose matrix mapping sequential process in Pose-Transformer with the aforementioned procedure. And figure 11 illustrate the structure of the Straw-Pose-Transformer.
@@
Figure 11: The architecture of the Straw-Pose-Transformer
@@
V combination of Pose-Transformer and baselines
@@ VI Experiment
VI-A Experiment on Bongard-Logo
We conducted experiments on the Bongard-logo dataset using the SBSD, PMoC and the integration of PMoC with Straw-Pose-Transformer. The results are recorded in Table I. Due to the serial convolution scanning mechanism and attention mechanism being prone to attention collapse, the parameters of the SBDC model are utilized as pre-trained parameters for the PMoC convolutional layer. It is worth noting that, under the experimental conditions of this paper, which involved four A100s graphics processing units (GPUs), the integration of PMoC with the Pose-Transformer proved to be challenging to deploy. This is one of the reasons and motivations behind our decision to design a lightweight approach for Pose-Transformer in this paper.
@@ VI-A Experiment on Bongard-Logo
TABLE I: Reasoning Accuracies of PMoC on Bongard-logo.
@@
PMoC 92.0 90.6 97.7 77.3 76.0
PMoC+Straw-Pose-Transformer 93.2 91.4 98.0 77.5 76.3
PMoC+Pose-Transformer Deployment is not feasible on our device
In Table I, it is evident that addressing the Bongard-Logo problem using a probabilistic model approach, namely PMoC, is both straightforward and effective. Furthermore, SBSD exhibits stronger outcomes when compared to previous models. Additionally, the integration of Straw-Pose-Transformer has contributed to enhancing the inference accuracy of PMoC.
Finally, we postulate that the significance of PMoC, a network primarily structured around the Transformer-Encoder, may be further emphasized when dealing with larger datasets. ViT [34] explicitly explores the importance of data volume for Transformer-based visual models, whereas RS-Tran [31] focuses on the significance of data quantity in addressing visual abstract reasoning challenges. Therefore, this paper attempts to conduct experiments on an expanded Bongard-Logo dataset to explore the further capabilities of PMoC and (Straw) Pose-Transformer. However, due to the limited availability of additional Bongard-Logo data for expanding the dataset, this paper uses the data augmentation techniques to achieve the objective. To expand the Bongard-Logo problems, we employ data augmentation techniques such as random rotating by integer angles ($90^{\circ}$, $180^{\circ}$, $270^{\circ}$) and random horizontal/vertical flipping. These augmentations are applied to implicitly increase the data volume without altering the essence of Bongard-Logo as a few-shot learning task or explicitly changing the level of reasoning difficulty. We anticipate that this augmentation approach will enable PMoC and (Straw) Pose-Transformer to unleash their greater potential. The results is illustrated in Table II.
TABLE II: Reasoning Accuracies of PMoC on Augmented Bongard-logo.
Accuracy(%)
Model Train FF BA CM NV
PMoC 94.5 92.6 98.0 78.3 76.5
PMoC+Straw-Pose-Transformer 98.7 96.3 98.3 79.5 76.6
PMoC+Pose-Transformer Deployment is not feasible on our device
This paper argues that the adopted data augmentation methods do not explicitly alter the core reasoning difficulty or logical complexity of the Bongard-Logo problem. These augmentation techniques, from an intuitive perspective, neither add nor remove any preexisting concepts. Even if these data augmentation methods implicitly introduce additional and unknown concepts to the Bongard-Logo dataset, it is important to note that such concepts are not present in the test set, as no data augmentation has been applied to it. Furthermore, they do not provide guidance for generalization issues such as NV and CM, as the solutions to these problems are not directly reflected in the HD problem after applying such augmentations. Additionally, the data augmentation approach employed in this study is uniformly applied to all 14 images within each Bongard-Logo instance, maintaining the nature of few-shot learning (i.e., without increasing the number of images within an instance) and without introducing additional priors (such as rotational and flipping invariance of individual images within Bongard-Logo instances). These augmentation methods indeed enrich the expressions of concepts at the pixel level, which is precisely the challenge we anticipate for PMoC through data augmentation and expanded data volume.
VI-B Experiment on RPM
We conducted experiments on I-RAVEN and PGM by sequentially replacing the Transformer structure in RS-Tran [31] with Pose-Transformer and Straw-Pose-Transformer. The results of I-RAVEN are recorded in Table III and PGM is recorded in Table IV. In this paper, we considered that compared to the standard Transformer-Encoder, a single layer of Pose-Transformer and its lightweight version, Straw-Pose-Transformer, possess twice the number of logical computations. Therefore, the combined RS-Pose-Tran and RS-Straw-Pose-Tran only stacked half of the logical layers of RS-Tran to ensure that the internal logical computation times of the three models were equivalent. In addition, the parameter settings and dataset configurations for the experiments related to RS-Pose-Tran were strictly maintained consistent with those of RS-Tran. Such a setting renders the comparative results of the three models more convincing.
TABLE III: Reasoning Accuracies on RAVEN and I-RAVEN.
@@ VI-B Experiment on RPM
In Table IV, it can be seen that we did not conduct experiments on RS-Pose-Tran for PGM. This is because compared to RS-Tran operating on the RAVEN dataset, the overall dimension of the RS-Tran model doubles on PGM. However, such a doubling is prohibitive for RS-Pose-Tran, making it undeployable on our equipment and device. This also motivated us to lightweight the Pose-Transformer into the Straw-Pose-Transformer in this paper.
TABLE IV: Reasoning Accuracies on PGM.
@@
RS-Straw-Pose-TRAN 98.6(98.64)
RS-Pose-TRAN Deployment is not feasible on our device
From the perspective of experimental results, it is not difficult to find that our lightweight approach, RS-Straw-Pose-Tran, has been successful without significantly reducing the performance of RS-Pose-Tran. Furthermore, the introduction of the pose mapping matrix into the Transformer-Encoder is effective, significantly improving the performance of RS-Tran. This demonstrates that Pose-Transformer is more suitable for abstract reasoning capabilities compared to the standard Transformer-Encoder, even though the standard Transformer-Encoder structure has shown strong advantages in RPM and Bongard-logo problems.
VII Conclusion
@@
In conclusion, our Pose-Transformer emphasizes the effectiveness of integrating the pose mapping matrix into the Transformer-Encoder framework, thereby elevating abstract reasoning capabilities. Furthermore, we have successfully deployed a streamlined variant of Pose-Transformer, termed Straw-Pose-Transformer, without compromising performance, which further reduces resource consumption in addressing complex reasoning challenges. Additionally, the innovative design of PMoC circumvents the need to directly constrain the latent representation distribution of Bongard-Logo images or rigidly model this distribution into predefined formats. Experimental results validate the rationality and effectiveness of the PMoC design approach. Consequently, PMoC, enhanced by Pose-Transformer, has exhibited its advantage in tackling the Bongard-Logo problem.